Wed Dec 31 11:41:14 PST 2003 — Web Server Logs: One of our webservers log rotation broke on the 26th. This lead to some problems running web stats and the replication of logs and stats for three days. Web logs and stats should be processed correctly tonight. -Kevan, Nathan and Kelsey
Wed Dec 31 16:07:53 PST 2003 — Media server change postponed. The switch to the new streaming media server has been postponed until Monday the 5th of January. This should be a transparent change for customers serving streaming media from their Sonic.net hosted web sites, and will not affect any other customers. -Kevan
Mon Dec 29 10:08:26 PST 2003 — Occidental Rooftop Network outage. Around 9:45am today the Occidental-area Rooftop Wireless network went offline. We are currently investigating the cause. No ETR available at this time, but this outage only affects customers with Rooftop Wireless in the Occidental area. – Bryan and John
Update: Outage was caused by a weather-related power outage. Service should be available again, power infrastructure permitting. – Bryan and John
Fri Dec 26 12:27:07 PST 2003 — Mail Server Upgrades: We’ve turned up a pair of new mail servers this morning to handle outbound email sent to mail.sonic.net. The servers are dual Xeons with 3-Ware SATA RAID using Western Digital Raptors. These boxes are blazing fast and will increase the reliability and performance of our outbound email flows. The two old servers will be taken offline next week, once their queued mail has been delivered or expired. -Kelsey and Nathan
Fri Dec 26 09:13:41 PST 2003 — Broadlink outage in Sector 3. All Broadlink Wireless connections served by the Barham Tower on Sector 3 are down this morning. Sonic.net Technical Support is coordinating with Broadlink Staff regarding this problem. There is no ETR at this time. – John and Eli
Update: Upon further review, it appears that this is not a sector-wide problem. A Broadlink monitoring tool issued information that appeared to be authoritative but was in fact errant. Customers on Sector 3 that are experiencing any connectivity issues are being handled as individual cases now. – John and Tim
Wed Dec 24 14:32:00 PST 2003 — Sonic is closing today at 2:30PM for the holidays. We will be reopening Friday morning at 8:00AM. Happy holidays from all of us at Sonic.
Mon Dec 22 21:41:42 PST 2003 — Mail Server NFS Filer Maintenance: This afternoon, one of the 2 pairs of clustered NFS filers that handle our message stores experienced some difficulty resulting in a cluster takeover by its partner. We don’t believe that any users experienced any impact when this occurred. Tonight, shortly after Midnight we will return the failed head to service. Unless something goes awry, it should take less than a minute to return it to service, during which time approximately half of our users may experience a delay checking their email. -Kelsey and Nathan
Sat Dec 20 16:24:12 PST 2003 — Basic Website Hosting product improved. We have increased the disk storage quota for our standard hosting product to 100 megabytes, plus added 24 POP e-mail boxes to the package.
This, plus the huge outbound bandwidth quota increases and inclusion of SQL database hosting which were made earlier this year have significantly increased the value and economy of our private-IP website hosting product (formerly known as “multihomed hosting”). -Dane
Sat Dec 20 16:19:17 PST 2003 — DSL outage in Oakland. SBC-ASI has informed us that they are currently experiencing an outage in a portion of Oakland which is affecting some Sonic.net customers. SBC-ASI is working now to resolve the issue. If you are in Oakland and are down, please dial up until the issue is resolved. -Dane and ScottB